### v1

@@ 3.1 On the $m$OneMinMax Problem
First, we prove that for any objective vector $\bm{v}$, the probability of not finding it in $O(m\mu n\log n)$ fitness evaluations is at most $n^{-m}$. For any solution $\bm{x}$, we partition it into $m/2$ blocks, where the $i$-th ($i\in[1..m/2]$) block $B_{i}=[(i-1)(2n/m+1)\ldots i(2n/m)]$. When there is a bit flip in block $\bm{x}_{B_{i}}$, one of $f_{2i-1}(\bm{x})$ and $f_{2i}(\bm{x})$ will increase by $1$, and the other one will decrease by $1$, such that $f_{2i-1}(\bm{x})+f_{2i}(\bm{x})=2n/m$ remains unchanged. Therefore, $\bm{x}$ needs to flip at least $\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ bits to obtain $\bm{v}$. Let $d_{\bm{v}}=\min_{\bm{x}\in P}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ denote the minimum number of bits that require to be flipped to obtain $\bm{v}$ for all the solutions in the population $P$. Since $d_{\bm{v}}=\min_{\bm{x}\in P}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2\leq\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2\leq\left\|2\bm{v}\right\|_{1}/2=n$, and $d_{\bm{v}}=0$ when there exists an $\bm{x}\in P$ such that $\bm{f}(\bm{x})=\bm{v}$, we have $0\leq d_{\bm{v}}\leq n$. Since all the objective vectors will be preserved, $d_{\bm{v}}$ cannot increase. Let the random variable $X_{k},k\in[1..n]$ denote the number of generations with $d_{\bm{v}}=k$. Let $X:=\sum_{k=1}^{n}X_{k}$. When $d_{\bm{v}}=k$, the probability of selecting an individual $\bm{y}$ from the population such that $\left\|\bm{f}(\bm{y})-\bm{v}\right\|_{1}/2=k$ is at least $1/\mu$. Since flipping any one of the $k$ bits corresponding to $d_{\bm{v}}$ will reduce $d_{\bm{v}}$ by $1$, the probability that $d_{\bm{v}}$ decreases is at least $(k/n)\cdot(1-1/n)^{n-1}\geq k/(en)$ when using bit-wise mutation and at least $k/n$ when using one-bit mutation. In the following, we consider bit-wise mutation, while the analysis for one-bit mutation holds analogously. In each generation, the probability that $d_{\bm{v}}$ decreases is at least:
@@
where $c\mu$ is the size of the offspring population, the second inequality holds by $1+a\leq e^{a}$ for any $a\in\mathbb{R}$, and the third inequality holds by $k\leq n/m$. Let $p_{k}:=ck/(c+e)n$. Then, $X_{k}$ stochastically dominates a geometric random variable $Y_{k}$ with success probability $p_{k}$. That is for any $\lambda\geq 0$, we have $\mathrm{P}(X_{k}\geq\lambda)\leq\mathrm{P}(Y_{k}\geq\lambda)$. Although $X_{1},\cdots,X_{n}$ are not independent, it can be guaranteed that the success probability of $X_{k}$ is at least $p_{k}$. Therefore, for $Y:=\sum_{k=1}^{n}Y_{k}$ where $Y_{1},\cdots Y_{n}$ are independent, $X$ stochastically dominates $Y$. By Theorem 16 in  [16] we have:
@@
$\mathrm{P}\Big{(}X\geq\frac{5en^{2}}{c}\Big{)}\leq\mathrm{P}\Big{(}Z\geq\frac{5en^{2}}{c}\Big{)}=\mathrm{P}(Z\geq E[Z]+\delta)\leq e^{-\frac{1}{4}\min\{\frac{\delta^{2}}{s},\frac{c\delta}{c+en}\}}\leq e^{-n}\,\text{,}$ (4)
where $\delta=5en^{2}/c-en\log n/c-n$, $E[Z]=en\log n/c+n$ and $s=\sum_{k=1}^{n}1/p_{k}^{\prime 2}=n+2en\log n/c+(e\pi n)^{2}/6c$. Thus, the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ after $c\mu\cdot 5en^{2}/c=5e\mu n^{2}$ fitness evaluations is at most $e^{-n}$.
@@ 3.2 On the $m$LeadingOnesTrailingZeroes Problem
We prove in Theorem 3.2 that expected the number of fitness evaluations for any MOEA solving $m$LeadingOnesTrailingZeroes is $O(\mu n^{2})$. The proof idea is to divide the optimization procedure into two phases, where the first phase aims at finding a point in the Pareto front. Let $w_{i}(\bm{x}):=f_{2i-1}(\bm{x})+f_{2i}(\bm{x})$ denote the sum of leading ones and trailing zeroes of block $\bm{x}_{B_{i}}$. Let $W_{\max}$ denote $\max_{\bm{x}\in P}\sum_{i=1}^{m/2}w_{i}(\bm{x})$. We will prove that the expected number of fitness evaluations for $W_{\max}$ to reach $n$ (i.e., a Pareto front point is found) is at most $O(\mu n^{2})$. The second phase aims at finding the whole Pareto front extended from the the Pareto front point found in the first phase. Similar to the analysis on $m$OneMinMax, we prove that the probability of not finding a specific point in the Pareto front within $6e\mu n^{2}$ number of fitness evaluations is at most $e^{-n}$ by applying Theorem 1 in [42]; by the union bound, the probability of finding the whole Pareto front within $O(\mu n^{2})$ number of fitness evaluations is at least $1-(2n/m+1)^{m/2}\cdot e^{-n}=1-o(1)$. Since the proof is similar to that of Theorem 4.3 in [36], and also resembles that of Theorem 3.1, we omit the detailed proof here.
@@ 3.3 On the $m$OneJumpZeroJump Problem
where the last inequality holds by $k\leq n^{\prime}/2$. The same bound also applies to $|\bm{x}_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}]$. Let $J(\bm{x})$ denote the number of blocks $B_{i}$ such that $|\bm{x}_{B_{i}}|_{1}\in[k..n^{\prime}-k]$, and let $J_{\max}:=\max_{\bm{x}\in P}J(\bm{x})$. For a solution $\bm{x}^{*}$ with $J(\bm{x}^{*})=J_{\max}$, it will not be dominated by any solution $\bm{y}$ with $J(\bm{y})<J_{\max}$. Thus, $J_{\max}$ will not decrease. When $J_{\max}$ reaches $m/2$, an internal Pareto front point is found. The probability of selecting the solution $\bm{x}^{*}$ corresponding to $J_{\max}$ is $1/\mu$, and the probability of increasing $J_{\max}$ by 1 through bit-wise mutation on $\bm{x}^{*}$ is at least $\frac{1}{e}(\frac{1}{mk})^{k}$ by Eq. (5). Therefore, in each generation, the probability of increasing $J_{\max}$ by $1$ is at least:
@@
The second phase is to find all internal Pareto front points. The analysis of the second phase is similar to that of $m$OneMinMax. After finding an internal Pareto front point, we show that the probability of finding all internal Pareto front points in $O(\mu n^{2})$ fitness evaluations is at least $1-o(1)$. For any internal Pareto front point $\bm{v}$, let $d_{\bm{v}}=\min_{\bm{x}\in F^{I}}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ denote the minimum number of bit flips required to obtain $\bm{v}$ for all internal Pareto front points found so far, the set of which is denoted as $F^{I}$. Since all internal Pareto front points will be preserved, $d_{\bm{v}}$ cannot increase. By Eq. (2), the probability that $d_{\bm{v}}$ decreases is at least $ck/(c+en)$. Let the random variable $X_{k}$ for $k\in[1..(m/2)(n^{\prime}-2k)]$ denote the number of generations with $d_{\bm{v}}=k$. Let $X:=\sum_{k=1}^{m(n^{\prime}-2k)/2}X_{k}$. Then similar to Eq. (3), we can derive that $\mathrm{P}(X\geq 5en^{2}/c)\leq e^{-n}$. In each generation, the population produces $c\mu$ offspring, thus the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ after $c\mu\cdot 5en^{2}/c=O(\mu n^{2})$ fitness evaluations is at most $e^{-n}$. For any internal Pareto optimum $\bm{x}$, $f_{2i-1}(\bm{x})+f_{2i}(\bm{x})=2k+n^{\prime}$, and the value range is $[2k,n^{\prime}]$. Hence, the size of the internal Pareto front is $(n^{\prime}-2k+1)^{m/2}$. By union bound, the probability of finding the whole internal Pareto front within $O(\mu n^{2})$ fitness evaluations is at least $1-(n^{\prime}-2k+1)^{m/2}\cdot e^{-n}=1-o(1)$.
@@
First, we show that the expected number of fitness evaluations for finding all the extreme points in the Pareto front is $O(3^{m/2}\cdot\mu n^{k})$. We divide the optimization procedure of the third phase into $m/2$ levels, and each solution at the $i$-th level contains $i$ extreme blocks. We first consider the running time of finding all the extreme Pareto front points at level $1$. After the second phase, all internal Pareto front points have been found. Hence, for any block $B_{i}$, there exist at least $\sigma_{1}:=(n^{\prime}-2k+1)^{m/2-1}$ solutions $\bm{x}$ such that $|\bm{x}_{B_{i}}|_{1}=k$. The probability of selecting such a solution as parent is at least $\sigma_{1}/\mu$, and the probability of flipping $k$ $1$-bits to obtain $\bm{y}$ such that $|\bm{y}_{B_{i}}|_{1}=0$ is $(1/n^{k})\cdot(1-1/n)^{n-k}\geq 1/(en^{k})$. The same result also applies for $|\bm{x}_{B_{i}}|_{1}=n^{\prime}-k$ to generate a solution $\bm{y}$ such that $|\bm{y}_{B_{i}}|_{1}=n^{\prime}$. Therefore, in each generation, the probability of generating an extreme block is at least $1-(1-\sigma_{1}/(e\mu n^{k}))^{c\mu}\geq 1-e^{-c\sigma_{1}/(en^{k})}\geq c\sigma_{1}/(c\sigma_{1}+en^{k})$, where the inequalities holds by $1+a\leq e^{a}$ for any $a\in\mathbb{R}$. Thus, the expected number of fitness evaluations for generating an extreme block is at most $c\mu\cdot(c\sigma_{1}+en^{k})/c\sigma_{1}=O(\mu n^{k}/\sigma_{1})$. Then, we consider finding all the solutions at level $1$ where $B_{i}$ is an extreme block. This optimization process is similar to the second phase, extending from an internal Pareto front point to the whole internal Pareto front, and the difference is no need to consider the block $B_{i}$. Therefore, the expected number of fitness evaluations of this process is less than the second phase, which is $O(\mu n^{2})$. Because there can be $m/2$ positions with extreme blocks, and each extreme block has $2$ possible forms $0^{n^{\prime}}$ and $1^{n^{\prime}}$, the expected number of fitness evaluations to find all level $1$ solutions is $O(m/2\cdot 2\cdot\mu n^{k}/\sigma_{1})+O(m/2\cdot 2\cdot\mu n^{2})=O(m\mu(n^{k}/\sigma_{1}+n^{2}))$. Next, we consider finding all solutions at level $2$. The number of combinations of $2$ extreme blocks from $m/2$ positions is $\binom{m/2}{2}\cdot 2^{2}$, and each combination can be obtained by adding an extreme block to the solution at level $1$. Assume that there is no solution $\bm{x}$ in the population such that both $B_{i}$ and $B_{j}$ (where $i\neq j$) are extreme blocks. After finding all solutions at level $1$, there exists at least $\sigma_{2}:=(n^{\prime}-2k+1)^{m/2-2}$ solutions that can be used to generate a solution at level $2$ (with $B_{i}$ and $B_{j}$ as the extreme blocks) by mutating $k$ bits. Then, following the above analysis, the expected number of fitness evaluations for generating a solution at level $2$ is at most $O(\mu n^{k}/\sigma_{2})$. The expected number of fitness evaluations to extend from a solution at level $2$ to all solutions at level $2$ is also less than that of the second phase, which is at most $O(\mu n^{2})$. Thus, the expected number of fitness evaluations to find all level $2$ solutions is $O(\binom{m/2}{2}\cdot 2^{2}\cdot\mu(n^{k}/\sigma_{2}+n^{2}))$. Continuing this way, for each level $i$, we have $\sigma_{i}:=(n^{\prime}-2k+1)^{m/2-i}$ and the expected number of fitness evaluations to find all Pareto front points from level $1$ to level $m/2$ is at most $O(\sum_{i=1}^{m/2}\binom{m/2}{i}\cdot 2^{i}\mu(n^{k}/\sigma_{i}+n^{2}))=O((2+1/(n^{\prime}-2k+1))^{m/2}\mu n^{k}+3^{m/2}\mu n^{2})$. Since $k\leq n^{\prime}/2$, the inequality $1/(n^{\prime}-2k+1)\leq 1$ holds, implying that the expected number of fitness evaluations for finding all the extreme Pareto front points is at most $O(3^{m/2}n^{k})$.
@@
Then we show that after the second phase, all the extreme Pareto front points can be found in $O(m\mu n^{k+1})$ number of fitness evaluations with probability $1-e^{-\Omega(mn)}$. After the second phase, all internal Pareto front points have been found. Assume that a Pareto front point $\bm{v}$ containing $d$ extreme values (the objective value is $k$ or $n^{\prime}+k$) has not been found, then there exists a solution $\bm{x}\in P$ such that except for the $d$ extreme values of $\bm{v}$, the remaining objective values are all equal to $\bm{v}$. Then, $\bm{v}$ could be extended from $\bm{x}$ by generating $d/2$ extreme blocks. Let the random variable $X_{k},1\leq k\leq m/2$, denote the number of generations to generate the $k$-th extreme block. Let $X:=\sum_{k=1}^{m/2}X_{k}$. By the analysis of extreme blocks, the probability of generating an extreme block in each generation is at least $c/(c+en^{k})$. Let $p_{k}:=c/(c+en^{k})$. Hence, $X_{1},\ldots,X_{m/2}$ stochastically dominates independent geometric random variables $Y_{1},\ldots,Y_{m/2}$, where the success probability of $Y_{k}$ is $p_{k}$. Moreover, for $Y:=\sum_{k=1}^{m/2}Y_{k}$, $X$ stochastically dominates $Y$. By Theorem 1 in  [42], we have
@@
where $\delta=mn^{k+1}/c-m(c+en^{k})/2c$, $E[Y]=m(c+en^{k})/2c$ and $s=m(c+en^{k})^{2}/2c^{2}$. Thus, the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ in $c\mu\cdot(mn^{k+1}/c)=m\mu n^{k+1}$ number of fitness evaluations is at most $e^{-\Omega(mn)}$. For Pareto optimal solutions, each block corresponds to two objective values, and the value range is $\{k,n^{\prime}+k\}\cup[2k..n^{\prime}]$. Thus, the size of the extreme Pareto front is $(n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2}$. By applying the union bound, the probability of finding all the extreme Pareto front points within $m\mu n^{k+1}$ number of fitness evaluations is $1-((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\cdot e^{-\Omega(mn)}\geq 1-n^{m/2}e^{-\Omega(mn)}=1-o(1)$. Let each $m\mu n^{k+1}$ fitness evaluations to find the Pareto front be an independent trial with success probability of $1-o(1)$. Thus, the expected number of fitness evaluations to find the Pareto front is at most $(1-o(1))^{-1}m\mu n^{k+1}=O(m\mu n^{k+1})$. Finally, combining the above analysis, the third phase needs $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$ number of fitness evaluations in expectation.
@@
Theorem 3.3 shows that for any MOEA preserving the non-dominated set, the expected number of fitness evaluations for solving $m$OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$. Note that when $m$ is a constant, the expected number of fitness evaluations is $O(\mu n^{k})$. Conversely, if the number $m$ of objectives is large, e.g., $m=n/4$, the expected number of fitness evaluations is $O(\mu n^{k+2})$.

### v2

@@ 3.1 On the $m$OneMinMax Problem
First, we prove that for any objective vector $\bm{v}$, the probability of not finding it in $O(m\mu n\log n)$ fitness evaluations is at most $n^{-m}$. For any solution $\bm{x}$, we partition it into $m/2$ blocks, where the $i$-th ($i\in[1..m/2]$) block $B_{i}=[(i-1)(2n/m)+1\ldots i(2n/m)]$. When there is a bit flip in block $\bm{x}_{B_{i}}$, one of $f_{2i-1}(\bm{x})$ and $f_{2i}(\bm{x})$ will increase by $1$, and the other one will decrease by $1$, such that $f_{2i-1}(\bm{x})+f_{2i}(\bm{x})=2n/m$ remains unchanged. Therefore, $\bm{x}$ needs to flip at least $\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ bits to obtain $\bm{v}$. Let $d_{\bm{v}}=\min_{\bm{x}\in P}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ denote the minimum number of bits that require to be flipped to obtain $\bm{v}$ for all the solutions in the population $P$. Since $d_{\bm{v}}=\min_{\bm{x}\in P}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2\leq\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2\leq\left\|2\bm{v}\right\|_{1}/2=n$, and $d_{\bm{v}}=0$ when there exists an $\bm{x}\in P$ such that $\bm{f}(\bm{x})=\bm{v}$, we have $0\leq d_{\bm{v}}\leq n$. Since all the objective vectors will be preserved, $d_{\bm{v}}$ cannot increase. Let the random variable $X_{k},k\in[1..n]$ denote the number of generations with $d_{\bm{v}}=k$. Let $X:=\sum_{k=1}^{n}X_{k}$. When $d_{\bm{v}}=k$, the probability of selecting an individual $\bm{y}$ from the population such that $\left\|\bm{f}(\bm{y})-\bm{v}\right\|_{1}/2=k$ is at least $1/\mu$. Since flipping any one of the $k$ bits corresponding to $d_{\bm{v}}$ will reduce $d_{\bm{v}}$ by $1$, the probability that $d_{\bm{v}}$ decreases is at least $(k/n)\cdot(1-1/n)^{n-1}\geq k/(en)$ when using bit-wise mutation and at least $k/n$ when using one-bit mutation. In the following, we consider bit-wise mutation, while the analysis for one-bit mutation holds analogously. In each generation, the probability that $d_{\bm{v}}$ decreases is at least:
@@
where $c\mu$ is the size of the offspring population, the second inequality holds by $1+a\leq e^{a}$ for any $a\in\mathbb{R}$, and the third inequality holds by $k\leq n$. Let $p_{k}:=ck/(c+e)n$. Then, $X_{k}$ stochastically dominates a geometric random variable $Y_{k}$ with success probability $p_{k}$. That is for any $\lambda\geq 0$, we have $\mathrm{P}(X_{k}\geq\lambda)\leq\mathrm{P}(Y_{k}\geq\lambda)$. Although $X_{1},\cdots,X_{n}$ are not independent, it can be guaranteed that the success probability of $X_{k}$ is at least $p_{k}$. Therefore, for $Y:=\sum_{k=1}^{n}Y_{k}$ where $Y_{1},\cdots Y_{n}$ are independent, $X$ stochastically dominates $Y$. By Theorem 16 in  [16] we have:
@@
$\mathrm{P}\Big{(}X\geq\frac{5en^{2}}{c}\Big{)}\leq\mathrm{P}\Big{(}Z\geq\frac{5en^{2}}{c}\Big{)}$ (4)
$=\mathrm{P}(Z\geq E[Z]+\delta)\leq e^{-\frac{1}{4}\min\{\frac{\delta^{2}}{s},\frac{c\delta}{c+en}\}}\leq e^{-n}\,\text{,}$
where $\delta=5en^{2}/c-enH_{n}/c-n$, $E[Z]=enH_{n}/c+n$ and $s=\sum_{k=1}^{n}1/p_{k}^{\prime 2}=n+2enH_{n}/c+(e\pi n)^{2}/6c$, with $H_{n}$ representing the $n$-th harmonic number, i.e. $\sum_{i=1}^{n}(1/i)$. Thus, the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ after $c\mu\cdot 5en^{2}/c=5e\mu n^{2}$ fitness evaluations is at most $e^{-n}$.
@@ 3.2 On the $m$LeadingOnesTrailingZeroes Problem
We prove in Theorem 3.2 that expected the number of fitness evaluations for any MOEA solving $m$LeadingOnesTrailingZeroes is $O(\mu n^{2})$. The proof idea is to divide the optimization procedure into two phases, where the first phase aims at finding a point in the Pareto front. Let $w_{i}(\bm{x}):=f_{2i-1}(\bm{x})+f_{2i}(\bm{x})$ denote the sum of leading ones and trailing zeroes of block $\bm{x}_{B_{i}}$. Let $W_{\max}$ denote $\max_{\bm{x}\in P}\sum_{i=1}^{m/2}w_{i}(\bm{x})$. We will prove that the expected number of fitness evaluations for $W_{\max}$ to reach $n$ (i.e., a Pareto front point is found) is at most $O(\mu n^{2})$. The second phase aims at finding the whole Pareto front extended from the Pareto front point found in the first phase. Similar to the analysis on $m$OneMinMax, we prove that the probability of not finding a specific point in the Pareto front within $6e\mu n^{2}$ number of fitness evaluations is at most $e^{-n}$ by applying Theorem 1 in [42]; by the union bound, the probability of finding the whole Pareto front within $O(\mu n^{2})$ number of fitness evaluations is at least $1-(2n/m+1)^{m/2}\cdot e^{-n}=1-o(1)$. Since the proof is similar to that of Theorem 4.3 in [36], and also resembles that of Theorem 3.1, we omit the detailed proof here.
@@ 3.3 On the $m$OneJumpZeroJump Problem
where the last inequality holds by $k\leq n^{\prime}/2$ (otherwise, internal Pareto optima will not exist). The same bound also applies to $|\bm{x}_{B_{i}}|_{1}\in[n^{\prime}-k+1..n^{\prime}]$. Let $J(\bm{x})$ denote the number of blocks $B_{i}$ such that $|\bm{x}_{B_{i}}|_{1}\in[k..n^{\prime}-k]$, and let $J_{\max}:=\max_{\bm{x}\in P}J(\bm{x})$. For a solution $\bm{x}^{*}$ with $J(\bm{x}^{*})=J_{\max}$, it will not be dominated by any solution $\bm{y}$ with $J(\bm{y})<J_{\max}$. Thus, $J_{\max}$ will not decrease. When $J_{\max}$ reaches $m/2$, an internal Pareto front point is found. The probability of selecting the solution $\bm{x}^{*}$ corresponding to $J_{\max}$ is $1/\mu$, and the probability of increasing $J_{\max}$ by 1 through bit-wise mutation on $\bm{x}^{*}$ is at least $\frac{1}{e}(\frac{1}{mk})^{k}$ by Eq. (5). Therefore, in each generation, the probability of increasing $J_{\max}$ by $1$ is at least:
@@
The second phase is to find all internal Pareto front points. The analysis of the second phase is similar to that of $m$OneMinMax. After finding an internal Pareto front point, we show that the probability of finding all internal Pareto front points in $O(\mu n^{2})$ fitness evaluations is at least $1-o(1)$. For any internal Pareto front point $\bm{v}$, let $d_{\bm{v}}=\min_{\bm{x}\in F^{I}}\left\|\bm{f}(\bm{x})-\bm{v}\right\|_{1}/2$ denote the minimum number of bit flips required to obtain $\bm{v}$ for all internal Pareto front points found so far, the set of which is denoted as $F^{I}$. Since all internal Pareto front points will be preserved, $d_{\bm{v}}$ cannot increase. By Eq. (2), the probability that $d_{\bm{v}}$ decreases is at least $cd_{\bm{v}}/(c+en)$. Let the random variable $X_{i}$ for $i\in[1..(m/2)(n^{\prime}-2k)]$ denote the number of generations with $d_{\bm{v}}=i$. Let $X:=\sum_{i=1}^{m(n^{\prime}-2k)/2}X_{i}$. Then similar to Eq. (4), we can derive that $\mathrm{P}(X\geq 5en^{2}/c)\leq e^{-n}$. In each generation, the population produces $c\mu$ offspring, thus the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ after $c\mu\cdot 5en^{2}/c=O(\mu n^{2})$ fitness evaluations is at most $e^{-n}$. For any internal Pareto optimum $\bm{x}$, $f_{2i-1}(\bm{x})+f_{2i}(\bm{x})=2k+n^{\prime}$, and $f_{2i-1}(\bm{x}),f_{2i}(\bm{x})\in[2k,n^{\prime}]$ for $i\in[1..m/2]$. Hence, the size of the internal Pareto front is $(n^{\prime}-2k+1)^{m/2}$. By union bound, the probability of finding the whole internal Pareto front within $O(\mu n^{2})$ fitness evaluations is at least $1-(n^{\prime}-2k+1)^{m/2}\cdot e^{-n}=1-o(1)$.
@@
First, we show that the expected number of fitness evaluations for finding all the extreme points in the Pareto front is $O(3^{m/2}\cdot\mu n^{k})$. We divide the optimization procedure of the third phase into $m/2$ levels, and each solution at the $i$-th level-contains $i$ extreme blocks. We first consider the running time of finding all the extreme Pareto front points at level-$1$. After the second phase, all internal Pareto front points have been found. Hence, for any block $B_{i}$, there exist at least $\sigma_{1}:=(n^{\prime}-2k+1)^{m/2-1}$ solutions $\bm{x}$ such that $|\bm{x}_{B_{i}}|_{1}=k$. The probability of selecting such a solution as parent is at least $\sigma_{1}/\mu$, and the probability of flipping $k$ $1$-bits to obtain $\bm{y}$ such that $|\bm{y}_{B_{i}}|_{1}=0$ is $(1/n^{k})\cdot(1-1/n)^{n-k}\geq 1/(en^{k})$. The same result also applies for $|\bm{x}_{B_{i}}|_{1}=n^{\prime}-k$ to generate a solution $\bm{y}$ such that $|\bm{y}_{B_{i}}|_{1}=n^{\prime}$. Therefore, in each generation, the probability of generating an extreme block is at least
$1-\Big{(}1-\frac{\sigma_{1}}{e\mu n^{k}}\Big{)}^{c\mu}\geq 1-e^{-\frac{c\sigma_{1}}{en^{k}}}\geq\frac{c\sigma_{1}}{c\sigma_{1}+en^{k}},$ (7)
where the inequalities hold by $1+a\leq e^{a}$ for any $a\in\mathbb{R}$. Thus, the expected number of fitness evaluations for generating an extreme block is at most $c\mu\cdot(c\sigma_{1}+en^{k})/(c\sigma_{1})=O(\mu n^{k}/\sigma_{1})$. Then, we consider finding all the solutions at level-$1$ where $B_{i}$ is an extreme block. This optimization process is similar to the second phase, extending from an internal Pareto front point to the whole internal Pareto front, and the difference is no need to consider the block $B_{i}$. Therefore, the expected number of fitness evaluations of this process is less than the second phase, which is $O(\mu n^{2})$. Because there can be $m/2$ positions with extreme blocks, and each extreme block has $2$ possible forms $0^{n^{\prime}}$ and $1^{n^{\prime}}$, the expected number of fitness evaluations to find all level-$1$ solutions is $O((m/2)\cdot 2\cdot\mu n^{k}/\sigma_{1})+O((m/2)\cdot 2\cdot\mu n^{2})=O(m\mu(n^{k}/\sigma_{1}+n^{2}))$. Next, we consider finding all solutions at level-$2$. The number of combinations of $2$ extreme blocks from $m/2$ positions is $\binom{m/2}{2}\cdot 2^{2}$, and each combination can be obtained by adding an extreme block to the solution at level-$1$. Assume that there is no solution $\bm{x}$ in the population such that both $B_{i}$ and $B_{j}$ (where $i\neq j$) are extreme blocks. After finding all solutions at level-$1$, there exists at least $\sigma_{2}:=(n^{\prime}-2k+1)^{m/2-2}$ solutions that can be used to generate a solution at level-$2$ (with $B_{i}$ and $B_{j}$ as the extreme blocks) by mutating $k$ bits. Then, following the above analysis, the expected number of fitness evaluations for generating a solution at level-$2$ is at most $O(\mu n^{k}/\sigma_{2})$. The expected number of fitness evaluations to extend from a solution at level-$2$ to all solutions at level-$2$ is also less than that of the second phase, which is at most $O(\mu n^{2})$. Thus, the expected number of fitness evaluations to find all level-$2$ solutions is $O(\binom{m/2}{2}\cdot 2^{2}\cdot\mu(n^{k}/\sigma_{2}+n^{2}))$. Continuing this way, for each level-$i$, we have $\sigma_{i}:=(n^{\prime}-2k+1)^{m/2-i}$ and the expected number of fitness evaluations to find all Pareto front points from level-$1$ to level-$m/2$ is at most $O(\sum_{i=1}^{m/2}\binom{m/2}{i}\cdot 2^{i}\mu(n^{k}/\sigma_{i}+n^{2}))=O((2+1/(n^{\prime}-2k+1))^{m/2}\mu n^{k}+3^{m/2}\mu n^{2})$. Since $k\leq n^{\prime}/2$, the inequality $1/(n^{\prime}-2k+1)\leq 1$ holds, implying that the expected number of fitness evaluations for finding all the extreme Pareto front points is at most $O(3^{m/2}\mu n^{k})$.
@@
Then we show that after the second phase, all the extreme Pareto front points can be found in $O(m\mu n^{k+1})$ number of fitness evaluations with probability $1-e^{-\Omega(mn)}$. After the second phase, all internal Pareto front points have been found. Assume that a Pareto front point $\bm{v}$ containing $d$ extreme values (the objective value is $k$ or $n^{\prime}+k$) has not been found, then there exists a solution $\bm{x}\in P$ such that except for the $d$ extreme values of $\bm{v}$, the remaining objective values are all equal to $\bm{v}$. Then, $\bm{v}$ could be extended from $\bm{x}$ by generating $d/2$ extreme blocks. Let the random variable $X_{i},1\leq i\leq m/2$, denote the number of generations to generate the $i$-th extreme block. Let $X:=\sum_{i=1}^{m/2}X_{i}$. By the analysis of extreme blocks, the probability of generating an extreme block in each generation is at least $c/(c+en^{k})$. Let $p_{i}:=c/(c+en^{k})$. Hence, $X_{1},\ldots,X_{m/2}$ stochastically dominates independent geometric random variables $Y_{1},\ldots,Y_{m/2}$, where the success probability of $Y_{i}$ is $p_{i}$. Moreover, for $Y:=\sum_{i=1}^{m/2}Y_{i}$, $X$ stochastically dominates $Y$. By Theorem 1 in [42], we have
@@
where $\delta=mn^{k+1}/c-m(c+en^{k})/(2c)$, $E[Y]=m(c+en^{k})/(2c)$ and $s=m(c+en^{k})^{2}/(2c^{2})$. Thus, the probability that the population contains no solution $\bm{x}$ with $\bm{f}(\bm{x})=\bm{v}$ in $c\mu\cdot(mn^{k+1}/c)=m\mu n^{k+1}$ number of fitness evaluations is at most $e^{-\Omega(mn)}$. For Pareto optimal solutions, each block corresponds to two objective values, and the value range is $\{k,n^{\prime}+k\}\cup[2k..n^{\prime}]$. Thus, the size of the extreme Pareto front is $(n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2}$. By applying the union bound, the probability of finding all the extreme Pareto front points within $m\mu n^{k+1}$ number of fitness evaluations is $1-((n^{\prime}-2k+3)^{m/2}-(n^{\prime}-2k+1)^{m/2})\cdot e^{-\Omega(mn)}\geq 1-n^{m/2}e^{-\Omega(mn)}=1-o(1)$. Let each $m\mu n^{k+1}$ fitness evaluations to find the Pareto front be an independent trial with success probability of $1-o(1)$. Thus, the expected number of fitness evaluations to find the Pareto front is at most $(1-o(1))^{-1}m\mu n^{k+1}=O(m\mu n^{k+1})$. Finally, combining the above analysis, the third phase needs $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$ number of fitness evaluations in expectation.
@@
Theorem 3.3 shows that for any MOEA preserving the non-dominated set, the number of fitness evaluations for solving $m$OneJumpZeroJump is $O(\mu n^{k}\cdot\min\{mn$, $3^{m/2}\})$ in expectation. Note that when $m$ is a constant, the expected number of fitness evaluations is $O(\mu n^{k})$. Conversely, if the number $m$ of objectives is large, e.g., $m=n/4$, the expected number of fitness evaluations is $O(\mu n^{k+2})$.